# On the Effects of Randomness on Stability of Learning with Limited Labelled Data: A Systematic Literature Review

### v1

@@ 2.3. Survey Methodology
(Bouthillier et al., 2021) $\times$
(Agarwal et al., 2021) $\times$ $\times$
(Boquet et al., 2019) $\times$ $\times$
(Lu et al., 2022) - $\times$ $\times$
(Zhao et al., 2021b) $\times$
(Osahor and Nasrabadi, 2022) $\times$ $\times$
(Zhang et al., 2021a) $\times$ $\times$
(Min et al., 2022) - $\times$ $\times$
(Chen et al., 2022) $\times$ $\times$
(Winata et al., 2021) - $\times$ $\times$
(Chang et al., 2021) $\times$ $\times$
(Schick and Schütze, 2021) - $\times$ $\times$
(Jundi and Lapesa, 2022) $\times$
(Wu et al., 2022) $\times$ $\times$
(Yang and Kwok, 2022) $\times$ $\times$
(Zhao et al., 2021a) - $\times$
(Setlur et al., 2021b) $\times$
(Dror et al., 2019) $\times$ $\times$
(Dodge et al., 2020) $\times$
(Setlur et al., 2021a) $\times$ $\times$
(Mosbach et al., 2021) $\times$ $\times$
(Kumar and Talukdar, 2021) - $\times$ $\times$
(Reimers and Gurevych, 2017) $\times$ $\times$
(Zhang et al., 2021b) $\times$
(Liu et al., 2022) - $\times$ $\times$
(Sellam et al., 2022) $\times$
(McCoy et al., 2020) $\times$
(Zhou et al., 2020) $\times$
(Bouthillier et al., 2019) $\times$ $\times$
(Dehghani et al., 2021) $\times$ $\times$
(Zhong et al., 2021) $\times$
(Dauphin and Schoenholz, 2019) $\times$ $\times$
(D’Amour et al., 2022) $\times$ $\times$
(Khurana et al., 2021) $\times$ $\times$
(Hua et al., 2021) $\times$ $\times$
(Hidey et al., 2022) $\times$ $\times$
(Pham et al., 2021) $\times$ $\times$
(Cioba et al., 2022) $\times$
(Phang et al., 2019) $\times$
(Wang et al., 2021a) $\times$ $\times$
(Zhang et al., 2022b) $\times$
(Lee et al., 2020) $\times$ $\times$
(Xu et al., 2021c) $\times$ $\times$
(Meng et al., 2022) $\times$ $\times$
(Zhang et al., 2022a) - $\times$
(Zheng et al., 2022) $\times$ $\times$
(Ye et al., 2021) $\times$
(Bragg et al., 2021) $\times$ $\times$
(Mukherjee et al., 2021) $\times$ $\times$
(Efrat et al., 2022) - $\times$ $\times$
@@ 3. Taxonomy for Literature Analysis and Categorisation
The identified core papers are categorised based on multiple dimensions that describe their characteristics of interest. We use two groups of characteristics: 1) primary categorisation characteristics, which provide the primary means for dividing and categorising the individual papers; and 2) secondary categorisation characteristics, which provide a more nuanced clustering of papers in the specific primary categorisation dimensions.
We define three separate primary properties that are used to categorise the papers: 1) what tasks are performed to address the randomness; 2) randomness factors addressed by the papers; and 3) what machine learning approach the papers use. The performed tasks property is the main property based on which we organise the rest of the survey, while the randomness factors and machine learning approach properties are mostly used to define the scope in this survey (see Section 2.1). We provide the paper distribution across individual randomness factors and specific machine learning approaches for all core papers in Table 2 (the paper distributions for specific tasks are included in the supplementary material, Section A).
Table 2. Relation between the randomness factors and machine learning approach across all analysed papers. We can observe different popularity of the randomness factors across different machine learning approaches. The majority focus being denoted in bold. For example, fine-tuning mostly focuses on the random seeds.
All Meta-Learning Fine-Tuning In-Context Learning
(N=8) (N=10) (N=24) (N=13)
Label Selection 1 2 2 1
Data Split 2 2 1
Task Choice 1 5 N/A
Data Choice 1 6 2 6
Random Seed 4 2 17 5
Model Initialisation 5 8 N/A
Order of Data 4 1 7 9
Implementation 1 1
Number of Samples 5 6
@@
Augmentation 1
Noise 1 1 1
Approach Specific N/A 2 3
@@ 3.1. Tasks Performed to Address the Randomness
This property specifies the different tasks for addressing the effects of randomness that are performed in the papers (the relations between these tasks, along with their input and output are illustrated in Figure 4). For this property, we define following 7 values:
@@
    Determine - an effort to discover the origin of the randomness (different from the randomness factors) is made by analysing the behaviour of the specific machine learning approach or model in presence of randomness. For example, to determine why changing the order of samples causes significant change in performance, the experiments that observe specific problems (e.g., label bias) are designed and analysed.
@@
    Report - it is proposed to report more than a single value when running and evaluating experiments and comparing approaches to better deal with the effects of randomness. Not necessarily dependent on experiments for addressing the effects of randomness, but also recommendations how to improve the reporting of results overall when the randomness has a strong effect on performance. For example, instead of reporting single average value, the minimum and maximum value over multiple runs should be reported as well.
@@
    Compare - a better comparison strategies between different approaches in presence of significant effects of randomness are proposed. For example, instead of comparing single average values, it is proposed to compare distributions of results using statistical tests.
@@
As the different papers can focus on different, often disjointed, tasks for addressing the effects of randomness, this property does not necessarily provide a distinct division of the papers. For example, the most common combination is to perform investigation, determination and mitigation of a specific randomness factor in a single paper. In such cases, the paper is mentioned in this survey multiple times, each time focusing only on the parts relevant for the specific task.
Figure 4. Different tasks for addressing the effects of randomness, along with their inputs, outputs and the relations between them. The dashed lines represent relations between the tasks that currently do not exist, but need to be considered in the optimal state of addressing the effects of randomness, e.g., using more sophisticated evaluation when determining the effectiveness of mitigation strategies. The related tasks are grouped together, e.g., evaluate task near the investigate task.
@@ 3.2. Randomness Factors
The randomness factors property specifies what randomness factors are actually addressed in the paper in some way (i.e., their effects are investigated or mitigated), but not factors that are just mentioned (e.g., mention that other papers found the factor to be unimportant). Papers can focus on multiple factors at the same time.
@@
The identified randomness factors can be categorised into three distinct groups: 1) factors related to the input data; 2) factors related to parameters of the model and the overall training process; and 3) factors related to the underlying implementation of deep learning frameworks and hardware.
As the randomness factors related to the input data, we identify following:
@@
    Label selection - specifies what data is considered to be annotated from larger set of unlabelled data. Represents the real world scenario with limited labelled budget, where only specific samples are annotated. Has potentially significant impact on the underlying data distribution.
@@
    Data split - randomness introduced by different splits of data into training, validation and testing set. Mainly impacts the information provided to model during training.
@@
    Augmentation - randomness introduced by different data augmentation strategies.
@@
    Noise - randomness introduced by adding noise to the different samples. In essence, it can be viewed as a subset of the augmentation randomness factor.
  8. (8)
    Number of samples - specifies the randomness introduced by choosing different number of labelled samples in training, such as different number of examples in in-context learning. It can be viewed as an edge case with regards to the definition of stability, as the expected change is often systematic (but is not always the case).
@@
    Random seed - randomness introduced by non-deterministic actions in the training process, such as the use of non-deterministic layers. Often also used as an aggregated factor for the parameter initialisation and the order of data randomness factors.
@@
    Hyperparameters - randomness caused by introducing small random changes into hyperparameters. Can also be used when performing hyperparameter optimisation in the low data regimes, due to large effects of randomness on them. Also aggregates the randomness introduced by changing parameters related to optimisation such as parameters used in optimiser or the number of training iterations.
@@
    Approach specific parameters - aggregates the randomness introduced by factors specific for only some of the approaches. Includes factors like different prompt design (in-context learning), or different task sampling strategies (meta-learning).
Randomness introduced by implementation of deep learning frameworks and hardware is a special low level group, which includes factors such as scheduling and floating point precision, auto-selection of primitive operations, how parallel processes are used, changes introduced by different implementation of frameworks such as TensorFlow or PyTorch and their different software versions, but also many other factors influencing reproducibility more than the stability of results (Pham et al., 2021; Gundersen et al., 2022). Since these randomness factors cannot be easily addressed, for the purpose of this survey, we do not specifically distinguish between them and consider them jointly as one group.
Each randomness factor is associated with a set of its randomness factor configurations (sometimes referenced in short as configurations). A single configuration specifies one state from the set of all possible states the randomness factor can attain. For example, a single randomness factor configuration for the order of data specifies one possible permutation of the samples, a single set of weights used for initialising the whole model, or a single random value to seed the random number generator. Using this definition, the set of randomness factor configuration can be infinite for some factors (e.g., initialisation weights) or finite, but extremely large (e.g., permutations of order).
The different randomness factors may depend on each other, leading to interactions between them. An example of such interactions is the relation between data choice and order of data randomness factors, where the choice of what data we use in training explicitly affects the order of samples. Therefore, such interactions can lead to negative, often unforeseen consequences, such as prohibiting reliable freezing of randomness factor configurations (e.g., it is not possible to have fixed order of samples when using different choices of data).

### v2

@@ 2.3. Survey Methodology
(Bouthillier et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$
(Agarwal et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Boquet et al., 2019) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Lu et al., 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Zhao et al., 2021b) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$
(Osahor and Nasrabadi, 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Min et al., 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Chang et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Schick and Schütze, 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Jundi and Lapesa, 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$
(Wu et al., 2022a) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Yang and Kwok, 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Zhao et al., 2021a) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$
(Setlur et al., 2021b) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\sqrt{}$
(Dror et al., 2019) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Dodge et al., 2020) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$
(Setlur et al., 2021a) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Mosbach et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Kumar and Talukdar, 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Reimers and Gurevych, 2017) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Zhang et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$
(Liu et al., 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Sellam et al., 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$
(McCoy et al., 2020) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\sqrt{}$
(Zhou et al., 2020) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\sqrt{}$
(Bouthillier et al., 2019) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Dehghani et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Zhong et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$
(Dauphin and Schoenholz, 2019) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(D’Amour et al., 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Khurana et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Hua et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Hidey et al., 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Pham et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Cioba et al., 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\sqrt{}$
(Phang et al., 2019) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\sqrt{}$
(Wang et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Zhang et al., 2022b) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\sqrt{}$
(Xu et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Meng et al., 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Zhang et al., 2022a) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$
(Zheng et al., 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Ye et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\sqrt{}$
(Bragg et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Mukherjee et al., 2021) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
(Efrat et al., 2022) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$ $\sqrt{}$$\times$ $\times$
@@ 3. Taxonomy for Literature Analysis and Categorisation
The identified core papers are categorised based on multiple dimensions that describe their characteristics of interest. We define the following three separate primary properties that are used to categorise the papers: 1) what tasks are performed to address the randomness; 2) randomness factors addressed by the papers; and 3) what machine learning (ML) approach the papers use. The performed tasks property is the main property based on which we organise the rest of the survey, while the randomness factors and machine learning approach properties are mostly used to define the scope in this survey (see Section 2.1). We provide the paper distribution across individual randomness factors and specific machine learning approaches for all core papers in Table 2 (the paper distributions for specific tasks are included in the digital appendix.
Table 2. Relation between the randomness factors and machine learning approaches across all analysed papers. We can observe different popularity of the randomness factors across different machine learning approaches. The majority focus is denoted in bold. For example, fine-tuning mostly focuses on the random seeds.
All Meta-Learning Fine-Tuning In-Context Learning Prompt-Based Learning PEFT
(N=13) (N=14) (N=35) (N=96) (N=17) (N=10)
Label Selection 1 3 4 3 1
Data Split 2 1 4 3 1
Task Choice 1 7 N/A
Data Choice 1 9 6 51 8
Random Seed 7 3 20 7 3 2
Model Initialisation 9 1 13 N/A 2 4
Order of Data 8 2 11 37 6 2
Implementation 5 1
Number of Samples 3 9 17 2 2
@@
Augmentation 3 1 1
Noise 1 2 1 1
Prompt N/A N/A N/A 46 11 5
Model Randomness 3 1 3 3 2 1
Approach Specific N/A 1 5 3
@@ 3.1. Tasks Performed to Address the Randomness
This property specifies the different tasks for addressing the effects of randomness that are performed in the papers (the relations between these tasks, along with their input and output are illustrated in Figure 4). For this property, we define the following 7 values:
@@
    Determine - an effort to discover the origin of the randomness (different from the randomness factors) is made by analysing the behaviour of the specific machine learning approach or model in the presence of randomness. For example, to determine why changing the order of samples causes a significant change in performance, the experiments that observe specific problems (e.g., label bias) are designed and analysed.
@@
    Report - it is proposed to report more than a single value when running and evaluating experiments and comparing approaches to better deal with the effects of randomness. Not necessarily dependent on experiments for addressing the effects of randomness, but also recommendations on how to improve the reporting of results overall when the randomness has a strong effect on performance. For example, instead of reporting a single average value, the minimum and maximum values over multiple runs should be reported as well.
@@
    Compare - better comparison strategies between different approaches in the presence of significant effects of randomness are proposed. For example, instead of comparing single average values, it is proposed to compare distributions of results using statistical tests.
@@
As the different papers can focus on different, often disjointed, tasks for addressing the effects of randomness, this property does not necessarily provide a distinct division of the papers. For example, the most common combination is to perform investigation, determination and mitigation of a specific randomness factor in a single paper. In such cases, the paper is mentioned in this survey multiple times, each time focusing only on the parts relevant to the specific task.
Figure 4. Different tasks for addressing the effects of randomness, along with their inputs, outputs and the relations between them. The dashed lines represent relations between the tasks that currently do not exist but need to be considered in the optimal state of addressing the effects of randomness, e.g., using more sophisticated evaluation when determining the effectiveness of mitigation strategies. The related tasks are grouped, e.g., evaluate task near the investigate task.
@@ 3.2. Randomness Factors
The randomness factors property specifies what randomness factors are addressed in the paper in some way (i.e., their effects are investigated or mitigated), but not factors that are just mentioned (e.g., mention that other papers found the factor to be unimportant). Papers can focus on multiple factors at the same time.
@@
The identified randomness factors can be categorised into three distinct groups: 1) factors related to the input data; 2) factors related to parameters of the model and the overall training process; 3) factors related to the underlying implementation of deep learning frameworks and hardware; and 4) factors that provide more systematic changes.
As the randomness factors related to the input data, we identify the following:
@@
    Label selection - specifies what data is considered to be annotated from a larger set of unlabelled data. Represents the real-world scenario with a limited labelled budget, where only specific samples are annotated. Has a potentially significant impact on the underlying data distribution.
@@
    Data split - randomness introduced by different splits of data into training, validation and testing set. Mainly impacts the information provided to the model during training.
@@
    Noise - randomness introduced by adding noise to the different samples.
@@
    Prompt format - randomness introduced by small changes in the prompt formats, such as replacing words with synonyms or words without semantic meaning.
@@
    Random seed - randomness introduced by non-deterministic actions in the training process. Often used as an aggregated factor for the parameter initialisation and the order of data randomness factors.
@@
    Model randomness - non-deterministic operations in the model, such as the use of non-deterministic layers or regularisation.
@@
    Approach specific parameters - aggregates the randomness introduced by factors specific for only some of the approaches (e.g., sampling strategy or number of iterations).
Randomness introduced by implementation of deep learning frameworks and hardware is a special low-level group, which includes factors such as scheduling and floating point precision, auto-selection of primitive operations, how parallel processes are used, changes introduced by different implementation of frameworks such as TensorFlow or PyTorch and their different software versions, but also many other factors influencing reproducibility more than the stability of results (Pham et al., 2021; Gundersen et al., 2022). Since these randomness factors cannot be easily addressed, for this survey, we do not specifically distinguish between them and consider them jointly as one group.
Finally, we also consider factors that provide more systematic changes and can be viewed as an edge case with regards to randomness. This group includes factors such as the randomness introduces by augmentation, hyperparameters or the number of samples. Besides these randomness factors, we also consider systematic choices as they can affect the sensitivity to the effects of randomness as well. These are often a result of the experimental setup and include choices such as the datasets and their characteristics (e.g., monolingual vs. multilingual), or the model size
Each randomness factor is associated with a set of its randomness factor configurations (sometimes referenced in short as configurations). A single configuration specifies one state from the set of all possible states the randomness factor can attain. For example, a single randomness factor configuration specifies one possible permutation of the samples (order), a single set of weights used for initialising the whole model, or a single random value to seed the random number generator. Using this definition, the set of randomness factor configuration can be infinite for some factors (e.g., initialisation weights) or finite, but extremely large (e.g., permutations of order).
The different randomness factors may depend on each other, leading to interactions between them (Pecher et al., 2024c; Weber et al., 2023; Chen et al., 2022a). An example of such interactions is the relation between data choice and order of data randomness factors, where the choice of what data we use in training explicitly affects the order of samples. Therefore, such interactions can lead to negative, often unforeseen consequences, such as prohibiting reliable freezing of randomness factor configurations (e.g., it is not possible to have a fixed order of samples when using different choices of data).